You have been managing a $5 million portfolio that has a beta of 1.05 and a required rate of return of 8.675%. The current risk-free rate is 5%. Assume that you receive another $500,000. If you invest the money in a stock with a beta of 0.85, what will be the required return on your $5.5 million portfolio? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

Currently:

8.675=5+1.05*(market rate-5)

(8.675-5)=1.05*(market rate-5)

market rate=(8.675-5)/1.05+5

=8.5%

Hence required rate for $500,000=5+0.85*(8.5-5)

=7.975%

Portfolio return=Respective return*Respective weight

=(5,000,000/5,500,000*8.675)+(500,000/5,500,000*7.975)

=8.61%(Approx).